Math 524  University of South Carolina
D. Meade  Fall 1995

Homework Problems for Chapter 4 Least Squares Optimization

Section Page Problems
4.1  152 # 1, 3, 7, 10, 11a
4.2  152 # 4, 8

\underline{Notes}

• Least squares problems are a special class of optimization problems. In general, the idea is to minimize the sum of the squares of several quantities. We will see that techniques studied earlier in this course can be applied to find the solution of these problems.

• Homework problems will be collected each Monday. In general, problems for each section are due on the first Monday following their discussion in class.

• Do not wait until the last minute to look at the problems. In fact, I suggest that you look at the problems in advance of our discussion of the material in class.

• I have tried to keep the number of problems to a minimum. However, some of these problems will require some thought. You should expect that some parts of them do not make sense the first time you read them. However, as you re-read the problem, and the text, you should find that the questions start to become clearer.

• I expect you to come to me for help whenever it’s needed. I believe it is very important (for this course) that you successfully work all these problems. My e-mail address is meade@math.sc.edu; my office phone is x6183.